A 30 mm steel rod when subjected to an axial tensile force was subjected to a strain of 0.60X10-3 . find the tensile force that caused the above strain. Take E = 200 KN/mm2 


1
Expert's answer
2021-04-16T07:47:03-0400

Strain on rod=0.6 ×10-3, E=. 200 kN/mm2, diameter of rod=30,radius ,r=15 mm

We know that

Stress= strain ×E

Stress= 0.6×10-3×200×1000 N/mm2

Stress= 120 N/mm2

Stress= "\frac{Tensile force}{area}"

"120 \u00d7 3.14 \u00d715\u00d715= T"

Tensile force= 84.78 kN